OpenStudy (anonymous):

HELP I WILL FAN AND MEDAL Nathan bought 8 wooden posts to build a fence. The length of each post is 6 feet 4 inches. What is the total length in inches of the 8 posts? A. 608 B. 576 C. 166 D. 76

OpenStudy (anonymous):

576

OpenStudy (yanasidlinskiy):

8 * 6 = 48 feet 48 * 12 = ________ inches 8 * 4 = _______ inches

OpenStudy (anonymous):

HOW DI YOU GET 576

OpenStudy (yanasidlinskiy):

@deshawn1 Show how you got it.

OpenStudy (anonymous):

8 * 6 = 48 feet 48 * 12 = ________ inches 8 * 4 = _______ inches

OpenStudy (neer2890):

6 feet 4 inches=72 inches +4 inches=76 inches so 1 wooden post =76 inches long and 8 wooden posts=76*8=608 inches

OpenStudy (anonymous):

i got 608 cause 8 * 6 = 48 feet 48 * 12 = ___576_____ inches 8 * 4 = __32_____ inches 576 +32 =608

OpenStudy (anonymous):

yea thats what i got i must have typed the wrong answer by mistake.